OpenStudy (anonymous):

x^2-6x+8=0 completing the square

OpenStudy (slaaibak):

x^2 - 6x + 8 + 1 - 1 = 0 x^2 - 6x + 9 -1 = 0 (x-3)^2 = 1

OpenStudy (anonymous):

x^2 - 6x + 8 + 1 - 1 = 0 x^2 - 6x + 9 -1 = 0 x^2 - 6x + 9 -1 = 0